l Review Careful measurements have been made of Olympic sprinters in the 100- meter dash. A quite realistic model is that the sprinters velocity is given by a(1-eR) where t is in s, vx is in m/s, and the constants a and b are characteristic of the sprinter. Sprinter Carl Lewiss run at the 1987 World Championships is modeled with a -11.81m/s and b -0.6887s ,

Part D Find an expression for the distance traveled at time t Express your answer in terms of the variables a, b, and t. AE -bt Submit Previous Answers Reguest Answer X Incorrect; Try Again; 9 attempts remaining Part E Your expression from part D is a transcendental equation, meaning that you cant solve it for t. However, its not hard to use trial and error to find the time needed to travel a specific perfect. distance. To the nearest 0.01 s, find the time Lewis needed to sprint 100.0 m. His official time was 0.01 s more than your answer, showing that this model is very good, but not
Answer #1

Part D :

We know that , velocity is defined as the time derivative of displacement. i.e.

  V) lt

So,

x = int v_{x} , dt

Given,

bt ur = a (1-e

And for Carl Lewis

a = 11.81 m/s

b = 0.6887 s-1

So,   a(1-

or, d -bt ..................................(1)

Where, C is integration constant which can be calculated using initial conditions.

The initial condition is : at time t = 0 second the position of the sprinters was at x = 0 m. Putting these values in equation (1), we get

  0 =a imes 0 + rac{a}{b} , e^{-b imes 0} + C

or,   dl 0

or,   C = -rac{a}{b}

Putting the value of C in equation (1), we get

x = at + rac{a}{b} , e^{-bt} - rac{a}{b} .........................................(2)

This is the required expression for distance travelled at time t.

Part E :

Putting the values of a and b in equation (2), we get

r= (11.81)?+ 11.81 -bt11.81 -bt є 0,6887 0.6887

or, r = (11.81)t + (17. 148) e-0.6887t-17.148

For x = 100 m , we can write

  100 = (11.81)t + (17.148) , e^{-0.6887t} - 17.148

or, 117.148= (11.81)t + (17.148) , e^{-0.6887t}

Dividing both sides by 11.81, we get

  9.919-t + 1.452e-0.6887 t ............................(3)

Now using trial and error method :

Let t = 10 seconds and put it into the RHS (right hand side) of equation (3) ,we get

R.H.S 101.452 e-0.6887xi

or,   = 10 + 1.452 imes 0.0010209 = 10.00148

Which is greater than the LHS (left hand side) of the equation (3) i.e. 9.919 .

let t = 9.9 second then

R.H.S= 9.9 + 1.452 , e^{-0.6887 imes 9.9} = 9.9 + 1.452 imes 0.00109 = 9.9015

Which is nearly equals to LHS of the equation (3).

Let t = 9.918 second then

  R.H.S 9.918 + 1.452e一0.6887×9.918ー9.91 1.452 × 0.001080 9.9195

Which is equals to the LHS of equation (3) .

Hence, the Lewis needed 9.918 second (or 9.92 second) to sprint 100 m .

For any doubt please comment and please give an up vote. Thank you.